### v1

# Commissioning of a radiofrequency quadrupole cooler-buncher for collinear laser spectroscopy

Abstract
A RadioFrequency Quadrupole (RFQ) cooler-buncher system has been developed and implemented in a collinear laser spectroscopy setup. This system is dedicated to convert a continuous ion beam into short bunches, while enhancing beam quality and reducing energy spread. The functionality of the RFQ cooler-buncher has been verified through offline tests with stable rubidium and indium beam, delivered from a surface ion source and a laser ablation ion source, respectively. With a transmission efficiency exceeding 60%, bunched ion beams with a full width at half maximum of approximately 2 $\mu$s in the time-of-flight spectrum have been successfully achieved. The implementation of RFQ cooler-buncher system has significantly improved the overall transmission efficiency of the collinear laser spectroscopy setup.
Radiofrequency quadrupole cooler-buncher, Collinear laser spectroscopy, Hyperfine structure, Time of flight
I Introduction
The fundamental properties of atomic nuclei, such as masses, nuclear spins, electromagnetic moments, and charge radii, are crucial for exploring exotic nuclear structures and probing the underlying nucleon-nucleon interactions Ye2025 ; YANG2023104005 ; YAMAGUCHI2021103882 . Laser spectroscopy techniques enable the precise determination of the four electromagnetic properties of ground and long-lived isomeric states of atomic nuclei by measuring the hyperfine structure (HFS) spectra of their atoms, ions, and even molecules. Collinear laser spectroscopy YANG2023104005 ; Blaum_2013 is one such technique capable of achieving high-resolution HFS spectrum measurements. This is realized by overlapping a fast ion beam ($\sim$30 keV) with lasers in either collinear or anti-collinear geometry, effectively suppressing the spectral Doppler broadening caused by the energy spread of the ion beam KAUFMAN1976309 .
There are two typical approaches used in CLS to measure the HFS spectrum. The most commonly used one is laser-induced fluorescence (LIF), which employs a continuous-wave narrow-band laser to excite atoms or ions from their ground or metastable states to higher excited states, followed by the detection of emitted fluorescence using photomultiplier tubes Nörtershäuser2010 . However, the experimental sensitivity of this approach is often limited by high background signals caused by stray laser light. This limitation could be tamed by delivering the ion beam in short bunches, which enables to gate data taking on the ion-bunch passage through the laser-beam interaction region. As a result, the signal-to-background ratio is improved by 3-4 orders of magnitude, as first demonstrated in Ref. PhysRevLett.88.094801 . Another approach for measuring the HFS spectrum using CLS is resonance ionization spectroscopy (RIS), which utilizes multiple lasers to stepwise excite and then ionize the targeted atoms VERNON2020384 . By detecting the resonantly laser-ionized ions with high efficiency, this approach eliminates the need for light detection, thereby further improving the overall sensitivity of CLS PhysRevC.96.041302 . In order to achieve high resonance ionization efficiency in this approach, high-power pulsed lasers are indispensable. Consequently, the ion beam must be delivered in a bunched mode to ensure proper temporal matching between the laser pulses and the ion beam pulses in the interaction region. An earlier attempt to use continuous ion beams for RIS in CLS resulted in limited efficiency due to the duty cycle losses ALKHAZOV1991400 . Therefore, a bunched ion beam is a precondition for high-resolution and high-sensitivity HFS spectrum measurements using CLS.
Figure 1: Hyperfine structure spectrum of 38K measured for $4s~{}^{2\!}S_{1/2}$to $4p~{}^{2\!}P_{1/2}$ atomic transition at BRIF-CIAE (a) WANG2022166622 and at ISOLDE-CERN (b) Koszorús2021 .
Figure 2: (a) Schematic of the RFQ cooler-buncher system, including the offline surface and laser ablation ion source, ion optics and the RFQ cooler-buncher. (b) Layout of overall control and HV system. See text for more details.
In recent years, we have been working on the development of a CLS setup using both LIF and RIS approaches at the Radioactive Ion Beam (RIB) facilities in China PKU-CLS-2022 ; WANG2022166622 . The first stage of the setup, based on the LIF approach, has already been implemented at the Beijing Radioactive Ion Facility (BRIF) of the CIAE. The first commissioning experiment successfully measured the HFS spectrum of the ${}^{38}\text{K}$ isotope, as shown in Fig. 1(a) WANG2022166622 . However, compared to the HFS spectrum (Fig. 1(b)) measured at CERN-ISOLDE using the Collinear Resonance Ionization Spectroscopy (CRIS) experiment Koszorús2021 , the achieved signal to background ratio in the HFS spectrum is relatively poor due to the use of a continuous ion beam. Furthermore, the obtained spectral resolution is significantly limited by large Doppler broadening ($\Gamma_{\rm D}=\nu_{0}{\delta E}/{\sqrt{2Emc^{2}}}$), which is caused by the considerable energy spread $(\delta E)$ of ion beam delivered from the BRIF facility WANG2022166622 .
The RadioFrequency Quadrupole (RFQ) cooler-buncher is a beam manipulation device designed to simultaneously satisfy the aforementioned beam requirements for CLS experiments, namely a pulsed ion beam with a low energy spread. This technique has been extensively implemented in the worldwide RIB facilities, such as IGISOL PhysRevLett.88.094801 , ISOLDE Mané:1198225 , NSCL MINAMISONO201385 , and TRIUMF BARQUEST2016207 ; Mané2011 , where it has proven effective in delivering high-quality ion beams for a range of applications.
In this work, we present a detailed commissioning test of a newly-installed RFQ cooler-buncher system, which was constructed based on a design kanitz2019construction ; MIRACLS_master by the MIRACLS collaboration at ISOLDE-CERN MAIER2023167927 . This system was tested by using a 30-keV stable Rb ion beam in continuous mode from surface ion source and a 30-keV stable In ion beam in bunched mode (with a typical bunch width of 100 $\mu$s) from a laser ablation ion source. Through a systemic test, we identified optimal operational parameters of RFQ cooler-buncher system for the upcoming CLS experiment. Under this condition, this system could provide an ion bunch with a 2-$\mu$s temporal width while maintain more than 60% overall transmission efficiency.
II RFQ cooler-buncher system
Figure 2(a) present a detailed schematic of the system, including the offline surface and laser ablation ion source, ion optics, the RFQ cooler-buncher, as well as its control and HV platform. Monovalent positive ions are produced by a surface or laser ablation ion source and accelerated to 30 keV PKU-CLS-2022 . The ion beam is reshaped by the electrostatic quadrupole triplet (QT1) lens for a better beam profile and transmission before injection into the RFQ cooler-buncher which is installed inside a six-way cross at a potential slightly lower than the 30-keV ion beam energy. Inside the cooler-buncher, the ions collide with the purified helium buffer gas to eventually reach thermal equilibrium with helium atoms. In this manner, any potential large energy spread of the incoming ion beam is reduced. Simultaneously, the combined radial RF and axial DC electric fields enable trapping and accumulation of ions, thereby converting the incoming continuous ion beam into an ion bunch with low (longitudinal and transverse) emittance, once released from the cooler-buncher. Following the ion extraction, the bunched ion beam is re-accelerated to 30 keV, and its beam profile is optimized by the QT optics, prior to being directed into the CLS beamline. Two Faraday cups (FCs) are installed upstream and downstream of the RFQ cooler-buncher to evaluate the ion beam transmission. The time structure, namely the time of flight (TOF) spectrum of the 30-keV bunched ion beam is recorded by a MagneTOF ion detector and data acquisition (DAQ) system CLS-DAQ-PKU .
II.1 Ion source
Two types of offline ion sources are used for the system test: a newly installed surface ion source that provides a stable ion beam in continuous mode, and a laser ablation ion source that offers a stable ion beam in bunched mode. The surface ion source, containing Rb atoms (HeatWave Labs, #101139), is mounted in a CF150 flange of a six-way cross as shown in Fig. 1 and adjusted with a linear actuator. This ion source could produce stable ion beams of ${}^{85}\text{Rb}$ (72.17%) and ${}^{87}\text{Rb}$ (27.83%) with an intensity of about 100 pA by using a 1.44-A heating current. It is worth noting that under these conditions, the average fluctuation of ion beam intensity remains within 1-pA for several hours, which is crucial for RFQ cooler-buncher test. Details of the laser ablation ion source can be found in Ref. PKU-CLS-2022 . In brief, a solid indium target is ablated with a 532-nm pulse laser, generating bunched stable ion beams of ${}^{113}\text{In}$ (4.28%) and ${}^{115}\text{In}$ (95.72%) with a temporal width of $\sim$ 100 $\mu$s.
II.2 Ion optics and beam diagnose
As shown in Fig. 2 (a), the stable ion beam, after being extracted from the ion source, was accelerated to 30 keV and monitored directly by the FC1. A voltage of -50 V is applied to a grating in front of the cup to suppress secondary electron emissions when the ion beam impinging the FC, ensuring accurate beam current readings at the level of 0.1 pA. A deflector installed after the FC is temporarily served as an equivalent beam gate. With an applied voltage controlled by a HV switch, this beam gate can be used to control the beam current injected into the subsequent RFQ cooler-buncher, keeping it below 1 pA. This is done to avoid potential space charge effects inside the RFQ cooler-buncher caused by the intense ion beam.
The QT1 lens upstream of the RFQ cooler-buncher is designed to optimize ion beam injection and ensure high efficient transmission during the subsequent injection and deceleration process. A grounding tube (GT1), 150 mm in length and with an inner diameter of 20 mm, is installed inside a ceramic insulator just after the QT1, which connects the beamline on ground potential with the floated vacuum chamber of the RFQ cooler-buncher. The purpose of GT1 is to provide a well-defined ground potential for the ions within the insulator tube, ensuring that the ions trajectory is controlled before injecting into the RFQ cooler-buncher. Additionally, this GT1 also functions as gas-flow restricting tube, separating the $10^{-7}\sim 10^{-6}$ mbar vacuum environment in the QT1 region and the $\sim 10^{-4}$ mbar vacuum inside the RFQ cooler-buncher sixway-cross (see Sec.II.3). The GT2, QT2, and FC2, located downstream of the RFQ cooler-buncher (Fig. 2(a)), have identical designs to those upstream, ensuring optimization of beam quality and evaluation of beam transport. All mentioned ion optics and FC are powered by $\pm$6-kV HV modules (EHS F0 60n/p) installed in a HV crate (ECH238, iseg).
Figure 3: (a) A schematic diagram depicting the internal structure of the RFQ cooler-buncher. (b) An illustration of the cooling and bunching mechanism. During the cooling process, the extraction endcap is at a high potential to trap the ions. When this high potential is switched off, the trapped ions are released in short bunches. (c) An illustration of the generated potential profile on the RF electrodes. (d) The voltage applied to the RF electrodes. It should be note that opposite electrodes have voltages with the same polarity, while adjacent electrodes carry voltages of the opposite polarity, which creates a periodic RF field that confines the radial motion of ions.
II.3 RFQ cooler-buncher
[…]
Upon entering the cooler-buncher, ions are radially confined by the RF quadrupole electric field (Fig. 3(c)) generated by RF electrodes, thereby ensuring optimal ion transmission along the axial direction and ion trapping with minimal loss. The RF electrodes are composed of four identical semi-cylindrical rods, each with a radius of 5 mm and a length of 152 mm, with a minimum separation of 20 mm between opposing rods, as shown in Fig.3(d). The RF power supply, designed and manufactured by Beijing BBEF Science & Technology Co., Ltd.(BGTPAX2231), provides two adjustable sine wave outputs with a frequency range of $f=0.3-1.5$ MHz and an amplitude of $U_{\rm{RF}}=0-200$V. These two identical RF signals, except for a 180 phase difference, are applied to the two pairs of opposing RF electrodes. Adjacent electrodes are driven 180 out of phase, while opposing electrodes share the same phase, thereby generating RF field in the cross-sectional plane of the RFQ electrodes, as illustrated in Fig.3(c). The effective confinement of the ions’ radial motion in the RF field is governed by the Mathieu equation major2005charged , which ensures stability when the parameter $0\leq q\leq 0.908$, defined as:
$q=\frac{eU_{\rm{RF}}}{m\pi^{2}f^{2}r_{0}^{2}}.$ (1)
Here, $r_{0}=10$ mm represents half the distance between the opposing RF electrodes.
The energy spread of ions can be reduced through collisions with buffer gas atoms inside the cooler-buncher during transmission. The selected gas element must have a high ionization potential to minimize the possibility of charge exchange during collisions, and a low atomic mass to guarantee the collisional cooling effect ISCOOL ; PhysRev.170.91 . In our system, helium with a purity of 99.999% was chosen as the buffer gas. The flow rate of the injected helium gas is adjusted by a flowmeter (Sevenstar, CS200) with a maximum capacity of 100 Standard Cubic Centimeter per Minute (SCCM), and the gas is introduced into the cooler-buncher via a gas feedthrough on the top flange of the cooler. As ions traverse the cooler-buncher, guided by the progressive electric field generating by the 12 DC electrodes (Fig. 3(a)), they also experiences continuous collisions with helium atoms. After a brief period, known as the cooling time - typically less than 1 ms, depending on ion mass and gas flow rate kanitz2019construction - ions eventually reach the thermal equilibrium with the buffer gas atoms.
Decelerated ions inside the cooler-buncher are also influenced by a DC electric field along the axial direction. This electric field, generated by 12 annular DC electrodes positioned along the axial direction (as shown in Fig.3(b)), guides the ion beam towards the end section. The first seven electrodes have a length of 13 mm, while the last five electrodes are 8 mm, allowing for more precise electric field control. The potentials of these 12 DC electrodes are supplied by ±500 V modules (iseg, EBS C0 05), forming an axial electric field together with the extraction endcap electrode, as shown in Fig.3(b). The design of the extraction electrodes mirror those of the injection section; however the extraction endcap is not directly connected to the HV module. Instead, it interfaces with the HV module via a Behlke Switch (GHTS 30), which is controlled by a periodic TTL signal. The Behlke Switch features two HV inputs (DC+ and DC-) from $\pm$6-kV HV modules (iseg, EHS 80 60n/p) and a single HV output, enabling the application of a periodic voltage to the extraction endcap. During a typical measurement cycle, the control TTL signal for the Behlke Switch remains at a low level, causing it to output a positive DC voltage (e.g. +1600 V), thereby maintaining a high potential at the end section, as shown in Fig.3(b). Under the combined effects of the RF field, collisions with buffer gas and DC electrostatic potential trap, ions are gradually cooled and accumulated at the end section. After a certain time of accumulation, the control TTL signal switches to a high level, the extraction endcap receives a negative voltage (e.g. -800 V) from Behlke Switch. Note that the voltage applied to the extraction cone is kept to a relative large negative value (e.g.-2000 V) for an efficient extraction. As a result, the ions trapped at the end section (Fig.3(b)) are released and re-accelerated to nearly 30 keV within a short time period of approximately $10~{}\mu$s, eventually forming an ion bunch with a temporal width of only a few microseconds.
To achieve the desired functionality of the RFQ cooler-buncher, multiple hardware components are required, including an RF power supply, buffer gas input, DC power supplies and others. All devices associated with this system are illustrated in Fig. 2(b). The entire cooler-buncher is floated at a high-voltage potential, HV1, supplied by a high-precision DC power supply (Heinzinger PNChp 40000-15pos). Consequently, all associated devices are installed on an HV platform at the same potential, HV1, within a HV cage, as shown in Fig. 2(b). A 50-kV isolation transformer is employed to electrically isolate HV1 from ground potential while supplying 220 VAC power to all the devices on the HV platform. To facilitate remote operation, all devices on the platform are controlled via pairs of photo-electric converters and optical fibers, ensuring bidirectional communication between the control PC and the equipment. This system allows for the control and/or monitoring of key components, such as the RF power supply, DC power supply, gas flowmeter, vacuum gauge controller and Behlke Switch. The TTL signals required to operate the Behlke Switch are generated by a pulse generator (Quantum Composer 9520).
III Performance Test
Figure 4: (a) Ion count rate and width of the time of flight spectrum of the 85Rb ion bunches from the RFQ cooler-buncher system as a function of the buffer gas flow rate. (b) Ion count rate of the 85Rb ion bunches as a function of the voltage amplitude of the applied RF signal at different RF frequencies.
Figure 5: The time of flight (TOF) spectra of stable 85Rb and 115In ions detected by the MagneTOF ion detector, with the RFQ cooler-buncher operation turned on and off. (a) TOF spectrum of 85Rb produced from the surface ion source. (b) TOF spectrum of 115Rb produced from the laser ablation ion source.
Table 1: A summary of the optimized operation parameters for the RFQ cooler-buncher system, tested with the Rb ions from the surface ion source. All values are directly read from device, except for DC potential gradient along the axial direction, which is obtained based on the COMSOL simulation using set voltages.
Parameter Value Parameter Value
Element Rb DC Potential Gradient (V/cm) 1.28
[…]
Buffer Gas Flowrate (SCCM) 20 RF Frequency (kHz) 500
The RFQ cooler-buncher system described above has been successfully manufactured and installed at Peking University. This system will be used to generate bunched ion beams for subsequent collinear resonance laser spectroscopy experiments. Thus, our primary focus is the TOF spectrum width of the ion bunches and its overall efficiency. In order to evaluate performance, the bunched ions are detected by a MagneTOF ion detector (ETP, 14924 MagneTOFTM Mini) located downstream in the laser spectroscopy setup, as shown in Fig.1 of Ref. CRIS-NIMB-PKU . The ion bunches generated from the cooler-buncher are delivered towards the ion detector through a 90 degree dipole magnet, as depicted in Fig. 1.
Given the large number of parameters involved in the RFQ cooler-buncher system operation, a systematic testing approach is required, including the testing of all the DC potentials, RF amplitude and frequency, buffer gas flowrate, and other factors. Prior to the experimental tests, simulation of this system was performed using COMSOL Multiphysics to identify suitable initial parameters for the testing process. Starting from these initial values, each component is tested and individually optimized, followed by iterative fine-tuning of various devices to achieve global optimization. For example, as shown in Fig. 4, the effects of the buffer gas flowrate and RF parameters on the ion bunches count rate and TOF width are systematically examined. Using stable Rb ions from the surface ion source, the maximum count rate of the ion bunches is achieved with an RF voltage of $U_{\rm{RF}}=167$ V, an RF frequency of $f=500$ kHz, and a helium flowrate of 20 SCCM, as shown in Fig. 4. Notably, throughout this optimization process, the width of TOF spectrum of ion bunches is maintained at approximately 2 $\mu$s, a value for this key parameter that is suitable for laser spectroscopy experiments.
Other parameters can also influence the properties of the bunched ion beam. For example, the DC- and DC+ voltages applied to the extraction endcap significantly affect both the ion count rate and the TOF spectrum. And the potential gradient across the 12 DC electrodes will change the profile of TOF spectrum (e.g., symmetrical vs. asymmetrical). In principle, a TOF spectrum with a width as narrow as few tens of nanoseconds can be achieved, as demonstrated at MIRACLS Carina2025 , but this comes at the cost of an increased energy spread and possibly a loss in overall efficiency. However, considering the commonly used TOF width of approximately 2 $\mu$s for collinear laser spectroscopy 79Zn ; CRIS-Fr , the optimized parameters for a transmission of 60% Rb ions are listed in Tab.1. Under these conditions, the observed TOF spectra of 85Rb ions from the surface ion source and 115In ions from a laser ablation ion source are shown in Fig.5. A comparison of the TOF spectra with and without RFQ cooler-buncher operation clearly demonstrates that this system effectively compresses the continuous ion beam and the bunched ion beam with a large temporal width of $\sim$100 $\mu$s, narrowing it into a significantly smaller ion bunches with a temporal width of $\sim$ 2 $\mu$s, while maintaining an overall efficiency greater than 60%.
Our ultimate goal is to integrate the RFQ cooler-buncher system into the newly-developed collinear resonance laser spectroscopy setup CRIS-NIMB-PKU , aimed at measuring the nuclear properties of unstable nuclei far from $\beta$-stability on the nuclear chart. However, as discussed in Ref. CRIS-NIMB-PKU , the earlier test of collinear resonance ionization laser spectroscopy, without the RFQ cooler-buncher system, encountered a major issue: an unacceptable overall ion beam transmission efficiency of less than 40%. This problem primarily arises from the distorted shape of the ion beam delivered into the CLS beamline, which significantly impacted the collinear overlap between the ion beam and laser beams within the meter-long interaction region, thereby reducing the resonance ionization spectroscopy efficiency PengZhangPhD . With the successful integration of the RFQ cooler-buncher between the ion source and the laser spectroscopy setup, we have now achieved an optimal overall ion beam transmission efficiency of $>$80%, from FC2 to the end section of the CLS beamline, as shown in Fig.1 of Ref. CRIS-NIMB-PKU . This demonstrates that, in addition to producing shorter ion bunches, the RFQ cooler-buncher system also improves the ion beam profile by reducing the beam’s transverse emittance.
IV Conclusion and Near term experiment plan
In summary, a RFQ cooler-buncher system and its commissioning test are presented. The results demonstrate that this system successfully fulfills its intended functions of compressing the ion beam into short bunches and improving the ion beam profile. With an efficiency of RFQ cooler-buncher exceeding 60%, ion beam bunches with a temporal width of $\sim$2 µs are achieved. Furthermore, the integration of this newly-developed system into the collinear laser spectroscopy setup has largely improved the overall transmission efficiency of the ion beam due to the enhanced beam profile.
Building on these commissioning results, the RFQ cooler-buncher system will be firstly used for the offline test of a collinear resonance ionization laser spectroscopy setup CRIS-NIMB-PKU and then for online laser spectroscopy experiment of unstable nuclei at BRIF facility BRIF targeting Rb and Cs isotopes. Using a UCX target, neutron-rich 85-100Rb and 147-150Cs isotopes have already been produced at BRIF facility with a yield exceeding 100 particles per second. Given the relatively large energy spread of the ion beam at BRIF, approximately 21 eV for 38K as observed in an earlier experiment WANG2022166622 , the integration of the RFQ cooler-buncher system will be crucial not only for producing bunched ion beams but also for reducing the energy spread of the radioactive ion beam. This will undoubtedly enhance the sensitivity and spectral resolution of the planned online collinear resonance ionization laser spectroscopy experiments.